Bob bought a goat.
He expected to sell it at a price that would give him a 10% profit on his purchase.
However, he had to sell it for $50 less than he expected, which was a loss of 15% on what it cost him.
What did Bob pay for the goat?

Answer by Earlsdon(6294)

Let C = the cost of the goat.
Expected sale price = C + 10% of C
Actual sale price = (C + 10% of C) - $ 50
Cost less 15% = C - 15% 0f C
C + 0.1C - $50 = C - 0.15C Simplify and solve for C
1.1C - $50 = 0.85C Add $50 to both sides.
1.1C = 0.85C + $50 Subtract 0.85C from both sides.
0.25C = $50 Divide both sides by 0.25
C = $200
Bob payed $200 for the goat.